OpenStudy (anonymous):

The base of triangle is 5cm greater than the height. The are a is 33cm^3. Find the height and the length of the base.

jhonyy9 (jhonyy9):

pardon but area not can being cm*3

OpenStudy (anonymous):

Given area of a triangle = 1/2(bh) and a base that is 5 greater than its height, we have b=h+5 (use this to get one variable) 33cm^2 = (1/2)(bh) or 33 = (1/2)(h(h+5) ---Remember the h+5 is still the base 33=(1/2)(h^2+5h) Multiply both sides by 2 to get 66=h^2+5h subtract 66 from both sides to get a quadratic equation: h^2+5h-66=0 Factor to (h+11)(h-6)=0; so h=-11 or h=6 since you can't have a negative height, pick h=6 Substitute b=h+5 to get b = 6+5 = 11 Answers: h=6, b=11 Check answers by inputting into A=(1/2)(bh) or A = (1/2)(66)=33